OpenStudy (anonymous):

Javier designs leather shoes at his workshop. He charges $45 for women's shoes and $50 for men's shoes. He needs to make $2,500 per week to meet his expenses. This situation can be given by 45x + 50y = 2,500, where x is the number of women's shoes and y is the number of men's shoes designed by Javier. If he does not have any female customers during a particular week, how many male customers does he need to meet his expenses? (x 0, y 0, x and y take only integer values)

OpenStudy (texaschic101):

45x + 50y = 2500 if no women's shoes(x) sold, then x = 0 now sub 0 in for x in the original equation 45(0) + 50y = 2500 50y = 2500 y = 2500/50 y = 50 --- he needs 50 male customers (that is, if they each buy 1 pair per person) so....(0,50)